# Space Group Informed Transformer for Crystalline Materials Generation

### v2

@@ IV.3 Plug-and-play materials design
Any property prediction model can used in conjunction with CrystalFormer for property-guided material generation. We train two crystal property prediction models using the MP-20 dataset and achieve mean absolute error (MAE) of 0.325 eV in band gap and MAE of 0.110 eV/atom in formation energy, which is comparable to the crystal graph convolutional neural network [87] trained on the same dataset [88]. We use the output of these two property prediction models as the forward probability $p(y|\boldsymbol{\mathcal{C}})$ of Eq. (3). More details are in appendix E.
Figure 9(a) demonstrates the controlled generation of materials with target band gap at $E_{\mathrm{g}}=2~{}\mathrm{eV}$ and the formation of energy $E_{\mathrm{form}}=-3~{}\mathrm{eV/atom}$ crystals [89, 90]. The conditional probability Eq. (3) contains both the model likelihood and the property regression MAE. Therefore, the generated samples will strike a balance between the two. To draw samples from the conditional probability distribution, we randomly generate a batch of 1000 crystal samples and sweep through the crystal sequence to update the atom species, fractional coordinates, and lattice parameters  [^2] For simplicity of the Wyckoff sequence are kept unchanged in the MCMC sampling.. Monte Carlo update of chemical species so to achieve the desired properties can be regarded as a systematic data-informed way of carrying out cation-transmutation for materials inverse design [92]. After reaching equilibrium, the histogram of band gap and formation energy is centered around the target values, which is shifted significantly away from the value of unconditionally generated samples. On the other hand, the likelihood shown in Fig. 9(b) indicates these conditional-generated crystals are not typical samples with respect to the unconditional distributuion. Neverthless, they are still probable samples according to the crystal prior given by the CrystalFormer. Given a myriad of materials property prediction models developed over the years and the inconvenience of re-training or fine-tuning the foundational generative model [16], we envision the plug-and-play generation approach demonstrated here to be a scalable way for materials design. We have exposed an interface of CrystalFormer in the code repository [34] for users to plug in arbitrary conditioners for guided materials generation.
V Related works
@@
The CrystalFormer is most closely related to the autoregressive generative model originally designed for molecules [99, 100, 101]. However, instead of predicting the relative distances of atoms, we predict the Wyckoff positions of symmetry-inequivalent atoms in the unit cell. Having the luxury of the space group symmetry for crystals provides strong hints on where to put the atoms in the unit cell and greatly simplifies the design around spatial symmetries. On the other hand, compared to Refs. [19] which treats text descriptions of crystals using autoregressive language models, we are only dealing with a more concise and essential atomistic representation of crystals, which leads to a smaller model size and faster sampling speed. Fast generation speed is not only a welcoming feature but also will be crucial for further exploration of materials space based on combinations of probabilistic generation and post-selection, Monte Carlo sampling, backtracking, and searching techniques [102]. More importantly, by baking in the space group symmetry in the model rather than learning them as statistical correlation from texts [18, 20], CrystalFormer guarantees space group constraints and cherishes the precious data and computing time. In a sense, the present work employs intrinsic mathematical (as opposed to natural) language to incorporate the symmetry principle in the generative modeling of crystals.
@@
VI Outlook
Precisely controlling the space group in the generative model of crystalline materials not only greatly simplifies the task but also is a highly desired feature for materials discovery and design. CrystalFormer integrates exact symmetry principles from math and empirical chemical intuitions from data into one unified framework. Probabilistic generative modeling of crystalline materials using CrystalFormer opens the way to many future innovations in materials discovery.
Note that the MP-20 dataset has by no means exhausted all available crystalline material [19, 16]. An obvious future direction is to scale up the model as well as the training dataset, especially curating a dataset with better coverage of space groups. In particular, extending the dataset to include both inorganic and organic crystals [108] may be beneficial as it improves the data coverage of low symmetric space groups. The transformer-based generative model is ready to be scaled up to work with much larger and more diverse training data, in the same fashion as large language models [109]. Given similar model architectures, the idea of generative pretraining of a foundational model for material generation is appealing. When scaling up the model it will be interesting to note the possible appearance of neural scaling law [110] as it has also been showing up in other contexts of atomistic modeling [111].
@@ VI Outlook
Conditioned materials generation depending on properties [101, 16, 21, 22] and experimental measurements [119] are highly desired features of materials generative model. Although it is straightforward to extend CrystalFormer (e.g. extend the space group embedding or employ the encoder-decoder transformer architecture [40]) to incorporate these conditions, we are particularly excited about the plug-and-play routine demonstrated in Sec. IV.3. Along this line, we envision an ecosystem [120] where the foundational generative model for $p(\boldsymbol{\mathcal{C}}|g)$ and more specialized discriminative models $p(y|\boldsymbol{\mathcal{C}})$ are developed separately but brought together via the Bayes rule.
@@ Acknowledgements.
We thank Han Wang, Lin Yao, Linfeng Zhang, Chen Fang, Yanchao Wang, Zhenyu Wang, Qi Yang, Shigang Ou, Xinyang Dong, Wenbing Huang, Quansheng Wu, Wanjian Yin, Xi Dai, Shuang Jia, Hangtian Zhu, Jiangang Guo, and Hongjian Zhao for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. T2225018, No. 92270107, No. 12188101, No. T2121001, and No. 12034009 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000 and No. XDB30000000.
@@ A.1 Model architectures
2:Parameters for the conditional probability of Wyckoff letters $\boldsymbol{\omega}_{i}$, chemical element $\boldsymbol{\alpha}_{i}$, and fractional coordinates $\boldsymbol{\chi}_{i},\boldsymbol{\upsilon}_{i},\boldsymbol{\zeta}_{i}$ of atoms and the lattice parameters $\boldsymbol{\ell}$.
3:$\boldsymbol{\omega}_{1}=\texttt{Net}(g)$triangleright the logit of the first Wyckoff position is implemented as a standalone neural network.
5:$\boldsymbol{h}_{\boldsymbol{W}}=[\texttt{Embed}(g),\texttt{Embed}(w_{i}),m_{i}]$.
6:$\boldsymbol{h}_{\boldsymbol{A}}=[\texttt{Embed}(g),\texttt{Embed}(a_{i})]$.
7:$\boldsymbol{h}_{\boldsymbol{X}}=[\texttt{Embed}(g),\cos(2\pi x_{i}),\sin(2\pi x_{i}),\ldots,\cos(2\pi x_{i}N_{f}),\sin(2\pi x_{i}N_{f})]$
8:$\boldsymbol{h}_{\boldsymbol{Y}}=\ldots$
9:$\boldsymbol{h}_{\boldsymbol{Z}}=\ldots$
@@
16:$\boldsymbol{\omega}_{i},\boldsymbol{\alpha}_{i},\boldsymbol{\chi}_{i},\boldsymbol{\upsilon}_{i},\boldsymbol{\zeta}_{i},\boldsymbol{\ell}=\texttt{Split}(\boldsymbol{h})$
@@
18:return $[\boldsymbol{\omega}_{1},\boldsymbol{\alpha}_{1},\boldsymbol{\chi}_{1},\boldsymbol{\upsilon}_{1},\boldsymbol{\zeta}_{1},\boldsymbol{\omega}_{2},\boldsymbol{\alpha}_{2},\boldsymbol{\chi}_{2},...,\boldsymbol{\ell}]$
@@
        Number of chemical species         119         ’H’ to ’Og’, plus padding atom
        Number of possible Wyckoff letters         28         ’a-z’+’A’, plus padding atom
@@
        Number of modes in lattice Gaussian mixture distribution $K_{l}$         16
@@ A.2 Sampling algorithm
1:space group number $g$, a list of chemical elements element_list, length $n$ of the atom sequence, sampling temperature $T$
@@
17:     Project ${x}$ to Wyckoff positions according to the Wyckoff letter $w$
@@
29:Get $\boldsymbol{\ell}$ from $\texttt{CrystalFormer}(g,\boldsymbol{W},\boldsymbol{A},\boldsymbol{X})$
30:$\boldsymbol{L}\sim\mathrm{GaussianMix}(\boldsymbol{\ell})^{1/T}$
@@
Algorithm 3 summarizes the Markov chain Monte Carlo sampling of CrystalFormer. It is used in the structure constrained generation of crystals in Sec. IV.2. The property-guided materials design discussed in Sec. IV.3 employs a similar sampling strategy for the posterior probability distribution Eq. (3).
@@
3:$\boldsymbol{\mathcal{C}}=(\boldsymbol{W},\boldsymbol{A},\boldsymbol{X},\boldsymbol{L})$
@@
17:Get $\boldsymbol{\ell}$ from $\texttt{CrystalFormer}(g,\boldsymbol{W},\boldsymbol{A},\boldsymbol{X})$
18:$\boldsymbol{L}\sim\mathrm{GaussianMix}(\boldsymbol{\ell})^{1/T}$
@@ Appendix B Validity and novelty of generated samples
Figure S2 illustrates the structure and compositional validity of generated samples across all 230 space groups. Following the Ref. [121], a structure meets the validity criteria if the shortest atomic distance exceeds $0.5~{}\AA$, a lenient standard. Composition validity requires charge neutrality as computed by SMACT [122]. This is, however, an overly stringent criterion since the composition validity of the training set is only around $90\%$ by this measure [11]. Note that the CrystalFormer can generate reasonable samples even for those space groups without any training data. This because the model can exploit knowledge learned from other space groups to place suitable atoms in the Wyckoff positions due to weight sharing. Moreover, since the sampling process makes use the of Wyckoff position table. The three dimensional coordinates of atoms are not completely random even for unseen space groups.
@@
[…]
@@
   PGSchNet [100] 99.65 75.96
   LM-CH (character-level tokenization) [18] 84.81 83.55
   LM-AC (atom coordinate-level tokenization) [18] 95.81 88.87
   Crystal-text-LLM [20] 96.5 86.3
@@
   CDVAE [11] 100.0 86.70
   DiffCSP [13] 100.0 83.25
   DiffCSP++ [26] 99.94 85.12
   UniMat-Large [15] 97.2 89.4
@@ Appendix C Discovered crystal samples with symmetric structure initialization
Table LABEL:table:samples and Table LABEL:table:pyxtal_samples list novel samples with $E_{\mathrm{hull}}<0.1$ eV/atom with CrystalFormer and PyXtal initializations respectively. Among them, Pd3Pt, MgAgPd2, TaNbRu2, Gd2HgAu, and TbMgPd2 are both discovered by CrystalFormer and PyXtal initialization. These compounds share identical crystal structures in the two tables, with lattice constants differing by less than 0.02 $\AA$.
@@
  • 1
    The fractional coordinates and lattice parameters are omitted for brevity. See details at https://drive.google.com/file/d/1gOIkWkjSH_Ed0-wzPk8VgxLjJcidbrH3/view?usp=sharing
@@
  • 1
    The fractional coordinates and lattice parameters are omitted for brevity. See details at https://drive.google.com/file/d/18sdsp-6yRSBaNez1A0lr20Ru6H7OOzvJ/view?usp=drive_link.
@@ Appendix D Discovered crystal samples with element substitution
Formula Space group Wyckoff-Atom sequence 1 $E_{\mathrm{hull}}$ (eV/atom) Lattice constant ($\AA$)
@@
Formula Space group Wyckoff-Atom sequence 1 $E_{\mathrm{hull}}$ (eV/atom) Lattice constant ($\AA$)
@@ Appendix E Details of plug-and-play materials design
We build two independent predictive models of band gap and formation energy following the generative pretraining strategy [128]. Generative pretrain of CrystalFormer allows the model to extract crystal representations. Then, by attaching a regression neural network to its feature and training on property labels, the model can be used to predict the properties of crystals.

### v3

@@ III.3 Plug-and-play materials design
Any property prediction model can be used in conjunction with CrystalFormer for property-guided material generation. We utilize two pre-trained MEGNet [50, 86] models to predict the band gap and formation energy, using the output of these two property prediction models as the forward probability $p(y|\boldsymbol{\mathcal{C}})$ of Eq. (3). More details are in appendix E.
Fig. 9a demonstrates the controlled generation of materials with target band gap at $E_{\mathrm{g}}=2~\mathrm{eV}$ and the formation of energy $E_{\mathrm{form}}=-3~\mathrm{eV/atom}$ crystals [87, 88]. The conditional probability Eq. (3) contains both the model likelihood and the property regression MAE. Therefore, the generated samples will strike a balance between the two. To draw samples from the conditional probability distribution, we randomly generate a batch of 1000 crystal samples and sweep through the crystal sequence to update the atom species, fractional coordinates, and lattice parameters  [^2] For simplicity of the Wyckoff sequence are kept unchanged in the MCMC sampling.. Achieving the desired properties via Monte Carlo update of chemical species can be regarded as a systematic data-informed way of carrying out cation-transmutation for materials inverse design [90]. After reaching equilibrium, the histogram of band gap and formation energy is centered around the target values, which is shifted significantly away from the value of unconditionally generated samples. On the other hand, the likelihood shown in Fig. 9b indicates these conditional-generated crystals are not typical samples with respect to the unconditional distribution. Nevertheless, they are still probable samples according to the crystal prior given by the CrystalFormer. Note that controlling the distribution of formation energy can significantly impact the distribution of the energy above the convex hull due to the correlation between these two energies. To assess the stability of the conditionally generated samples, we first use M3GNet [74] to filter out the unstable materials, followed by DFT verification on the remaining subset. The conditionally generated samples with $E_{\mathrm{hull}}<0.1$ eV/atom that are not included in the MP-20 dataset are listed in Table LABEL:table:plug_and_play_samples in appendix E. We observed that some materials meet the requirements of property predictors, but not the DFT calculations. This is due to due to the errors in the property prediction models, which could be further improved by employing a more robust crystal property prediction models. Given a myriad of materials property prediction models developed over the years and the inconvenience of re-training or fine-tuning the foundational generative model [16], we envision the plug-and-play generation approach demonstrated here to be a scalable way for materials design. We have exposed an interface of CrystalFormer in our code repository for users to plug in arbitrary conditioners for guided materials generation.
IV Discussion and Conclusion
IV.1 Related works
@@
The CrystalFormer is most closely related to the autoregressive generative model originally designed for molecules [97, 98, 99]. However, instead of predicting the relative distances of atoms, we predict the Wyckoff positions of symmetry-inequivalent atoms in the unit cell. Having the luxury of the space group symmetry for crystals provides strong hints on where to put the atoms in the unit cell and greatly simplifies the design around spatial symmetries. On the other hand, compared to Ref. [19] which treats text descriptions of crystals using autoregressive nature language model, CrystalFormer speaks native crystal language: it deals with a more concise and essential atomistic representation of crystals, which leads to a smaller model size and faster sampling speed. Fast generation speed is not only a welcoming feature but also will be crucial for further exploration of materials space based on combinations of probabilistic generation and post-selection, Monte Carlo sampling, backtracking, and searching techniques [100]. More importantly, by baking in the space group symmetry in the model rather than learning them as statistical correlation from texts [18, 20], CrystalFormer guarantees space group constraints and cherishes the precious data and computing time. In this sense, the present work employs rigorous mathematical (as opposed to vague natural) language to incorporate the symmetry principle in the generative modeling of crystals.
@@
IV.2 Outlook
Precisely controlling the space group in the generative model of crystalline materials not only greatly simplifies the task but also is a highly desired feature for materials discovery and design. CrystalFormer integrates exact symmetry principles from math and empirical chemical intuitions from data into one unified framework. Probabilistic generative modeling of crystalline materials using CrystalFormer opens the way to many future innovations in materials design and discovery.
Note that the MP-20 dataset has by no means exhausted all available crystalline material [19, 16]. An obvious future direction is to scale up the model as well as the training dataset, especially curating a dataset with better coverage of space groups. A later version of CrystalFormer is trained on curated Alex-20 dataset [106, 107] has shown significantly improved performance [108]. In particular, extending the dataset to include both inorganic and organic crystals [109] may be beneficial as it improves the data coverage of low symmetric space groups. The transformer-based generative model is ready to be scaled up to work with much larger and more diverse training data, in the same fashion as large language models [110]. Given similar model architectures, the idea of generative pretraining of a foundational model for material generation is appealing. When scaling up the model it will be interesting to note the possible appearance of neural scaling law [111] as it has also been showing up in other contexts of atomistic modeling [112].
@@ IV.2 Outlook
Conditioned materials generation depending on properties [99, 16, 21, 22] and experimental measurements [120] are highly desired features of materials generative model. Although it is straightforward to extend CrystalFormer (e.g. extend the space group embedding or employ the encoder-decoder transformer architecture [39]) to incorporate these conditions, we are particularly excited about the plug-and-play routine demonstrated in Sec. III.3. Along this line, we envision an ecosystem [121] where the foundational generative model for $p(\boldsymbol{\mathcal{C}}|g)$ and more specialized discriminative models for materials properties $p(y|\boldsymbol{\mathcal{C}})$ are developed separately but brought together via the Bayes rule.
Conflict of Interest
The authors declare that they have no conflict of interest.
@@ Acknowledgements.
We thank Han Wang, Lin Yao, Linfeng Zhang, Chen Fang, Yanchao Wang, Zhenyu Wang, Qi Yang, Shigang Ou, Xinyang Dong, Wenbing Huang, Quansheng Wu, Wanjian Yin, Xi Dai, Shuang Jia, Hangtian Zhu, Jiangang Guo, and Hongjian Zhao for useful discussions. This project is supported by the National Natural Science Foundation of China (T2225018, 92270107, 12188101, T2121001, and 12034009), National Key Projects for Research and Development of China (2021YFA1400400), and the Strategic Priority Research Program of Chinese Academy of Sciences (XDB0500000 and XDB30000000).
Author contributions
Zhendong Cao and Lei Wang wrote the code and trained the model. Xiaoshan Luo performed the numerical calculations. Jian Lv and Lei Wang designed the research and supervised the project. All authors contributed to the data analysis, discussion of the results, and preparation of the manuscript.
Data availability
We have released the codes and trained model at https://github.com/deepmodeling/CrystalFormer.
@@ A.1 Model architectures
[…]
@@
[…]
@@
[…]
@@
   Number of chemical species    119    ‘H’ to ‘Og’, plus padding atom
   Number of possible Wyckoff letters    28    ‘a-z’+‘A’, plus padding atom
@@
   Number of modes in lattice Gaussian mixture distribution $K_{l}$    4
@@ A.2 Sampling algorithm
1:space group number $g$, a list of chemical elements element_list, maximum number of Wyckoff positions $n$, sampling temperature $T$
@@
17:  Project x{x} to Wyckoff positions according to the Wyckoff letter $w$
@@
[…]
@@
Algorithm 3 summarizes the Markov chain Monte Carlo sampling of CrystalFormer. It is used in the structure constrained generation of crystals in Sec. III.2. The property-guided materials design discussed in Sec. III.3 employs a similar sampling strategy for the posterior probability distribution Eq. (3).
@@
[…]
@@
[…]
@@ Appendix B Validity and novelty of generated samples
[…]
@@
Space group Crystal system Training samples Validity ($\%$) $\uparrow$ Params
@@
   PGSchNet [98] 99.65 75.96 -
   LM-CH (character-level tokenization) [18] 84.81 83.55 $\sim1$$0$ M
   LM-AC (atom coordinate-level tokenization) [18] 95.81 88.87 $\sim1$$0$ M
   Crystal-text-LLM [20] 96.5 86.3 70 B
@@
   CDVAE [11] 100.0 86.70 4.5 M
   DiffCSP [13] 100.0 83.25 12.3 M
   DiffCSP++ [26] 99.94 85.12 12.3 M
   UniMat-Large [15] 97.2 89.4 -
@@ Appendix C Discovered crystal samples with symmetric structure initialization
[…]
@@
Figure S3: (a)-(g) The distribution of per-atom volume changes before and after relaxation in seven representative space groups.
@@ Appendix D Discovered crystal samples with element substitution
[…]
@@
[…]
@@ Appendix E Details of plug-and-play materials design
We use two pre-trained MEGNet models [50, 86] for band gap and formation energy prediction, which is implemented in the MatGL [129] repository. It is important to note that we do not train the models from scratch or fine-tune them on the MP20 dataset. Instead, we utilize the pre-trained models provided in the repository. This approach closely aligns with actual practices in materials science, where numerous high-quality machine learning property predictors are readily accessible.
@@
Table LABEL:table:plug_and_play_samples shows a list of novel samples close to the energy hull with formation energy and band predicted by the MEGNet [50, 86] and DFT calculations.
Table S7: Samples discovered by plug-and-play manner with $E_{\mathrm{hull}}<0.1$ eV/atom which are not in the MP-20 dataset. The energy above the convex hull is computed via DFT while the formation energy and band gap are predicted both by DFT and machine learning model.
Formula Space group Wyckoff-Atom sequence $E_{\mathrm{hull}}$ (eV/atom) Formation energy (eV/atom) Band gap (eV)
DFT MEGNet DFT MEGNet
LiMoF6 225 a-Mo-b-Li-e-F -0.4816 -3.2893 -2.8811 0.0039 2.3289
Sr2ZrWO6 225 a-Zr-b-W-c-Sr-e-O -0.4642 -3.6203 -3.0082 0.0030 1.6535
Sr2HfWO6 225 a-Hf-b-W-c-Sr-e-O -0.4222 -3.6537 -3.1408 0.0126 1.9169
K2NaWF6 225 a-W-b-Na-c-K-e-F -0.2964 -3.1146 -2.8592 1.2693 1.8388
Cs2RbMoF6 225 a-Mo-b-Rb-c-Cs-e-F -0.2905 -3.1418 -2.9865 2.2306 1.8929
KCrF6 225 a-K-b-Cr-e-F -0.2802 -2.9034 -2.8221 0.0006 2.3546
Ba2YMnO6 225 a-Y-b-Mn-c-Ba-e-O -0.2439 -3.1016 -2.7970 0.0028 2.0245
Ba2ZrMnO6 225 a-Zr-b-Mn-c-Ba-e-O -0.2082 -3.2689 -3.0276 0.9285 1.8680
Sr2HfMnO6 225 a-Hf-b-Mn-c-Sr-e-O -0.1649 -3.3213 -3.0912 0.8717 1.8699
Ba2LaVO6 225 a-V-b-La-c-Ba-e-O -0.1624 -3.3785 -3.0244 1.3322 1.5306
Cs2RbCrF6 225 a-Rb-b-Cr-c-Cs-e-F -0.1562 -3.1799 -2.9984 2.2985 3.2548
Cs2RbVF6 225 a-V-b-Rb-c-Cs-e-F -0.1356 -3.2064 -3.1524 0.0014 1.7538
Cs2NaVF6 225 a-Na-b-V-c-Cs-e-F -0.1104 -3.2037 -3.1418 0.0004 2.0226
Cs2MnTlF6 225 a-Mn-b-Tl-c-Cs-e-F -0.0666 -2.8449 -2.8082 0.0010 2.2036
K2MnCoF6 225 a-Mn-b-Co-c-K-e-F -0.0513 -2.8679 -2.7951 0.0038 2.1415
K2RbVF6 225 a-V-b-Rb-c-K-e-F -0.0159 -3.1498 -3.0416 0.0010 1.5841
HoMgPt2 225 a-Mg-b-Ho-c-Pt 0.0212 -0.9985 -1.0616 0.0215 -0.0074
TmPt 221 a-Pt-b-Tm 0.0563 -1.2660 -1.2784 0.1493 -0.0091
K2SrCoF6 225 a-Sr-b-Co-c-K-e-F 0.0706 -3.0990 -3.1450 0.0018 1.1738
HoTmIr2 225 a-Tm-b-Ho-c-Ir 0.0712 -0.8510 -0.8649 0.0125 -0.0074
Li2VF6 225 a-V-c-Li-e-F 0.0729 -3.0782 -2.9730 0.0015 1.8764
Lu2IrPt 225 a-Pt-b-Ir-c-Lu 0.0788 -1.0960 -1.1430 0.0047 -0.0076
Cs2LiCrF6 225 a-Cr-b-Li-c-Cs-e-F 0.0843 -3.0344 -2.9774 1.3975 1.9886
ErLuPt2 225 a-Lu-b-Er-c-Pt 0.0850 -1.2383 -1.2755 0.0066 -0.0098
HoLuPt2 225 a-Lu-b-Ho-c-Pt 0.0944 -1.2223 -1.2622 0.0327 -0.0102
ErPt 221 a-Pt-b-Er 0.0951 -1.2140 -1.2607 0.0691 -0.0089